OpenStudy (anonymous):

Solve the simultaneous equation: x+y=6 2x+y=8 I don't want a graph on this one... Show me how to use the addition method and/or the substitution method. Sorry I got brain freeze on this one x_x

OpenStudy (cwrw238):

since each equation contains +y, if we subtract the y will disappear x+y=6 2x+y=8 subtract: ------- -x = -2 x = 2 Now substitute x = 2 into first equation: 2 + y = 6 y = 6-2 = 4 solution x=2,y=4 thats it - done

OpenStudy (anonymous):

the answer is (2,3) .... so it's suppose to be x=2, y=3 ....

OpenStudy (cwrw238):

no thats wrong try plugging in x=2 and y=4 into each of the equations: 2 + 4 = 6 and 2(2) + 4 = 8

OpenStudy (cwrw238):

I used the elimination message some teachers call it the addition method and would multiply the first equation by -1 so the equations are then added to eliminate y, this is not really necessary as you can just subtract

OpenStudy (anonymous):

Nvm, I found the real solution. And it is (2,3) :-) Thank you for taking your time though.